BEP, ROE, AND ROIC Broward Manufacturing recently reported the following information: Net income ROA Interest expense Accounts payable and accruals $1,050,000 Browards tax rate is 35%. Broward finances with only debt and common equity, so it has no preferred stock, 40% of its total invested capital is debt, while 60% of its total invested capital is common equity. Calculate its basic earning power (BEP), its return on equity (ROE), and its return on invested capital (ROIC). Round your answers to two decimal places. BEP ROE ROIC $445,000 7% $151,300
Answer #1

Earnings before tax (EBT) = Net income / ( 1 - Tax % ) = $445,000 / ( 1 - 35% ) = $684,615

Earnings before interest and tax (EBIT) = EBT + Interest expenses = $684,615 + $151,300 = $835,915

ROA = Net income / Total assets

7% = $445,000 / Total assets

Total assets = $445,000 / 7% = $6,357,143

Total invested capital = Total assets - Account payables and accruals = $6,357,143 - $1,050,000 = $5,307,143

Common equity = 60% * Total invested capital = 60% * $5,307,143 = $3,184,286

Basic earning power (BEP) = EBIT / Total assets = $835,915 / $6,357,143 = 13.15%

Return on equity (ROE) = Net income / Equity = $445,000 / $3,184,286 = 13.97%

Return on invested capital (ROIC) = EBIT ( 1 - tax% ) / Total invested capital = $835,915 ( 1 - 35% ) / $5,307,143 = 10.24%